OpenStudy (anonymous):

Find the range of each function when the domain is {-2,1,7} 1.f(x)=5-4x 2.f(x)=-1/2x -8

OpenStudy (anonymous):

@lgbasallote @Callisto

OpenStudy (anonymous):

help! please

OpenStudy (anonymous):

replace \(x\) by \(-2\) then by 1, then by 7

OpenStudy (anonymous):

your three answers will be the range. for example if \(x=-2\) then \(.f(x)=5-4x=f(-2)=5-4\times (-2)=5+8=13\)

OpenStudy (anonymous):

repeat with 1 and 7, list those answers

OpenStudy (anonymous):

so the range would be.... {13,1,-23)?

OpenStudy (anonymous):

@satellite73

OpenStudy (phi):

For 1.f(x)=5-4x you want to find f(1) and f(7) (sat figured out f(-2) is 13) to do this, replace x with 1 in 5-4x (remember 4x is short hand for 4*x (4 times x)) 5-4*1= 5-4=1 5-4*7= 5-28= -23 so it looks like you got it.